OpenStudy (anonymous):

Write the equation of a line with a slope of 3/4 and a y-intercept of -3?

OpenStudy (anonymous):

possible answers: y = 3/4x - 3 y = 3/4x + 3 y = 3x - 3/4 y = -3x + 3/4

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

y=3/4 x-3

OpenStudy (anonymous):

thanks! can you help me with one more please?!

OpenStudy (anonymous):

yes sure:)

OpenStudy (anonymous):

:) Write an equation of a line with slope 4 and contains the point (0, 5)?

OpenStudy (anonymous):

possible answers y = 4x - 5 y = 5x - 4 y = 4x + 5 y = 5x + 4

OpenStudy (anonymous):

here., m=4 so, equation would be y=4x+b now we plug the point (0,5) in this and find the value of "b"

OpenStudy (anonymous):

so, can you guess the equation?

OpenStudy (anonymous):

y= 4x + 5

OpenStudy (anonymous):

yes correct:)

OpenStudy (anonymous):

wuuw! awesome thanks!

OpenStudy (anonymous):

my pleasure:)